A radioactive material A (decay constant ΠA) decays into a material B (decay constant ΠB) and then into C(decay constant ΠC) which is also radioactive. Determine the amount of C remaining after a time t.
1
Expert's answer
2020-05-28T11:48:45-0400

"A (ПА)\to B(ПА)\to C(ПА)"


"N_A(t)=N_{A0}e^{-ПА\cdot t}"


"N_{В0}=N_{A0}-N_A(t)=N_{A0}-N_{A0}e^{-ПА\cdot t}=N_{A0}(1-e^{-ПА\cdot t})"


"N_В(t)=N_{B0}e^{-ПВ\cdot t}=N_{A0}(1-e^{-ПА\cdot t})\cdot e^{-ПВ\cdot t}"


"N_С(t)=N_{С0}e^{-ПС\cdot t}"


"N_{С0}=N_{В0}-N_В(t)=N_{A0}(1-e^{-ПА\cdot t})-N_{A0}(1-e^{-ПА\cdot t})e^{-ПВ\cdot t}="


"=N_{A0}(1-e^{-ПА\cdot t})\cdot (1-e^{-ПВ\cdot t})"


"N_С(t)=N_{С0}e^{-ПС\cdot t}=N_{A0}(1-e^{-ПА\cdot t})\cdot (1-e^{-ПВ\cdot t})e^{-ПС\cdot t}"


"\frac{N_С(t)}{N_{A0}}=(1-e^{-ПА\cdot t})\cdot (1-e^{-ПВ\cdot t})e^{-ПС\cdot t}" Answer